OpenStudy (anonymous):

if, a*b=48 and a^2+b^2=100, find value of a and b?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Mk, looking at the first problem, do you know how you would solve for a?

OpenStudy (anonymous):

a = 48/b a + b = 10 if you take the square root of both sides of the second equation

OpenStudy (anonymous):

Now you have 2 equations and 2 unknowns

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

simaltaneously solve them

OpenStudy (anonymous):

a+b=10 and a*b=48?

OpenStudy (anonymous):

a = 48/b

OpenStudy (anonymous):

from there, substitute a = 48/b into a+b=10

OpenStudy (anonymous):

Which will be 48/b + b = 10

OpenStudy (anonymous):

but how does one get a+b=10 from a(square)+b(square)=100

OpenStudy (zarkon):

if, a*b=48 and a^2+b^2=100, find value of a and b? \[a=\frac{48}{b}\] \[\left(\frac{48}{b}\right)^2+b^2=100\] solve for b

OpenStudy (zarkon):

I get b=-8,-6,6,8

OpenStudy (anonymous):

solving it, gives (2304/b^2)+b^2=100,, right?

OpenStudy (anonymous):

solving it, we get b^2-10b+48=0 !!!!

OpenStudy (anonymous):

is this right?

OpenStudy (zarkon):

no